OpenStudy (anonymous):

When the polynomial P(x) = 5x3 − 51x2 + kx − 9 is divided by x − 9, the remainder is 0. Which of the following is also a factor of P(x)?

OpenStudy (anonymous):

A. x − 5 B. x + 1 C. x − 1 D. 5x + 1

OpenStudy (campbell_st):

well the easy solution is to use the factor theorem and ingore any polynomial of synthetic division so the 1st task is to find the value of k... so if x - 9 is a factor then x = 9 is a zero so substitute x = 9 into the polynomial to find the value of k... so \[0 = 5(9)^3 -51(9)^2 + 9k -9\] then when you get k... look at the other factors... the constant term is -9 so if (x - 9) is a factor... then (x -5) can't be... it would give a constant term that is at least 45 so look at x + 1 if its a factor then x = -1 is a possible zero same for x -1 is its a factor x = 1 is a possible zero... and the repeat it for 5x + 1 then x = 1/5 is a possible zero

OpenStudy (anonymous):

can you help with another question please

OpenStudy (campbell_st):

just post the question as a new question and lots of people can help